OpenStudy (anonymous):

solve by clearing fractions first. must show work. 1/2+3x=3/2x-5/2

OpenStudy (anonymous):

1/2+5/2=3/2x-3x 3=-3/2x x=-1/3

OpenStudy (anonymous):

what about multiplying every term by the lcd? how do i fit that in?

OpenStudy (anonymous):

help??!!!!

OpenStudy (anonymous):

wait, what is lcd? english is not my native language

OpenStudy (anonymous):

its ok:) but i got part of it hang tite please